Solve by substitution or elimination method:
3x-2y=26
-7x+3y=-49

Answer by roberta022469(1)

3(3x-2y)=26(3)
2(-7x+3y)=-49(3)
9x-6y=78
-14x+6y=-147
-------------
-5X=-69
x=13.8